OpenStudy (anonymous):

A motor bike travels 684 km in 9n hours going upstream and 378 km in three hours going down stream, What is the rate of the current?

OpenStudy (anonymous):

ok gimme a sec

OpenStudy (anonymous):

ehhh i dont know sorry

OpenStudy (kropot72):

"A motor bike travels 684 km in 9n hours going upstream and 378 km in three hours going down stream, What is the rate of the current?" 1. Should "....in 9n hours....." be written ".......in 9 hours......" ? 2. Is the motor bike adapted for travel on water?

OpenStudy (kropot72):

Let the bike speed in still water = b km/h Let the rate of the current = r km/h Speed going upstream = b - r = 684/9 Speed going downstream = b + r = 378/3 So now we have a pair of equations as follows: b - r = 76 ...........(1) b + r = 126 ........(2) If you add equations (1) and (2) you will eliminate r and can solve to find b. Then substitute the value of b into equation (1) and solve to find r.

OpenStudy (kropot72):

@learning1Are you there?